Supposefour coins are tossed. Let X represents the number of tails that occur. Illustrate a probability distribution of a random variable X USING FRACTION


1
Expert's answer
2021-05-02T08:15:31-0400

There are "2^4=16" outcomes.


"HHHH, HHHT, HHTH, HTHH,"

"HHTT, HTHT, HTTH,HTTT,"

"TTTT, TTTH, TTHT, THTT,"

"TTHH, THTH, THHT, THHH"

Let "X" represents the number of tails that occur.



"\begin{matrix}
 x & 0 & 1 & & 3 & 4 \\
\\
 p(x) & \dfrac{1}{16} & \dfrac{4}{16} & \dfrac{6}{16} & \dfrac{4}{16} & \dfrac{1}{16}
\end{matrix}"


"mean=E(X)=\sum_ix_ip(x_i)"

"=0(\dfrac{1}{16})+1(\dfrac{4}{16})+2(\dfrac{6}{16})+3(\dfrac{4}{16})+4(\dfrac{1}{16})=2"



"E(X^2)=\sum_ix_i^2p(x_i)"

"=0^2(\dfrac{1}{16})+1^2(\dfrac{4}{16})+2^2(\dfrac{6}{16})+3^2(\dfrac{4}{16})+4^2(\dfrac{1}{16})=5"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2=5-(2)^2=1"


"\sigma=\sqrt{\sigma^2}=1"